Solve by the substitution method.
4x+4y=-4
-2x+y=11
What is the solution to the system?
Help with this too please?!?!!

Answer by scott8148(6628)

adding 2x to the 2nd equation __ y=11+2x

SUBSTITUTING into the 1st equation __ 4x+4(11+2x)=-4 __ 4x+44+8x=-4 __ 12x=-48 __ x=-4

SUBSTITUTING into the 2nd equation __ -2(-4)+y=11 __ 8+y=11 __ y=3